wiki:UC20171113

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Phonecon, 11/13/17

Attending: Nancy Chanover (NMSU), Rachel Kuzio de Naray (GSU), Mark Klaene (APO), Joanne Hughes Clark (Seattle U), Ben Williams (UW), Mary Beth Kaiser (JHU), Scott Anderson (UW), Jon Holtzman (NMSU), Russet McMillan (APO; minute taker), Sarah Tuttle (UW), John Wilson (UVA), Bill Ketzeback (APO)


User feedback and comments from institutional representatives:

  • Colorado - no report
  • GSU - need both sides of DIS (red and blue) - one is not more important than the other
  • JHU - nothing to report
  • NAPG - no report
  • NMSU - good student training trip last weekend
  • Oklahoma - no report
  • Seattle - nothing to report
  • UVA - nothing to report
  • Washington - nothing to report
  • Wyoming - no report

Telescope and Instruments Report:

Mark's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 10/04/17 through 11/08/17

0) Overview

Reasonably good weather this period. ARCSAT observing. DIS and Echelle scattered light being monitored closely, both holding steady but at values worse than we thought were good. DIS blue camera vacuum serviced during a lull.

1) Telescope

Telescope was operational with minimal issues.

2) Instruments

DIS: With a short lull in the schedule it was decided to warm and pump on the blue camera. This is now complete and did reset the scatter light values we see. However the trend is still generally worse than we like. DCam spare was installed and the main camera is being sent to Andor to address bias level issue.

TripleSpec is operational.

Agile has been operational this period except for about a 2 week time when it was off schedule and the ICC died.

Echelle is operational. Inter-Order Light is higher than we like but improving slightly.

NIC-FPS is operational.

ARCTIC is operational with no change.

3) ARCSAT - General operations have been nominal but loss of throughput has been seen in data. Plans are in work to address this in Quarter 1 2018 with polishing and recoating mirrors. SurveyCam is still at Andor with no ETA.


Additional telescope and instrument discussion:

The echelle inter-order light is improving, which is expected as we move into the colder and drier winter months. DIS has been having some ion pump controller problems. Agile was down for a few day; its ICC computer has been fixed. We're not hearing much from Andor about SurveyCam and when we can expect it to be repaired and shipped back from Andor; we will update users when we learn more.


Q4 3.5-m scheduling:

There is some OPEN and DD05 time still available, all in December. Users should get their requests in soon if interested.


Q4 ARCSAT scheduling:

Requests received for listed OPEN time, but if anyone is interested, ask soon.


Q1 3.5-m scheduling:

Allocations have all gone out; now awaiting responses.


Q1 ARCSAT scheduling:

Because the primary mirror will be sent out for re-polishing in Q1, there are only a few weeks for which the telescope will be scheduled for science observations. There will be no official call for proposals; users should contact Ben and Russet soon if interested.


Debrief from ARC Board of Governors Meeting

Those from this call who attended the BoG meeting included Nancy, Scott, Mark, Jon H, and Mike S. The agenda was packed so there was not much time to convey a lot of information. Nancy gave her Director's Annual Report presentation (thank you to everyone who contributed science highlights!) and is considering how to streamline this process in future, for example by going over less of the information contained in the presentation and reserving more time for the discussion of a few key issues. Nancy highlighted a few items in the presentation, namely the activities of TDAWG, the status of instrument upgrades, and her plans for strategic planning. The Board had a few questions and comments concerning interest in the new spectrograph capabilities and thoughts on strategic coupling between SDSS and 3.5m (particularly with the new funding received from the Sloan Foundation for SDSS-V).


TDAWG Update

The Time Domain Astronomy Working Group (TDAWG) has been meeting regularly through biweekly telecons, compiling a website with lots of information about science use cases and potential recommendations for the 3.5m management (see https://sites.google.com/a/uw.edu/tdawg/home). Strategies for dealing with ToOs are still under discussion. There have been several recent high-profile ToOs, including the merging neutron stars in August and observations of the interstellar asteroid, which made the news recently.


TripleSpec Slit Width

Mike Skrustskie has been reluctant to allow slit changes, and recommends it should only be done with great care. The mechanism appears to be working normally but if it fails that could take the instrument down for an extended period until a servicing trip by the UVA team could be planned. Nancy would like to know how strong the science drivers are for changing slits at all. Action item for all UC reps: Please notify users of constraints on Tspec slit use and forward any interest in using other slits to Nancy.


NA2 Baffle Project

John Wilson and his team are making progress with the SolidWorks design of the physical structure of telescope and a Zmax model of optics. ARCTIC flats in multiple filters are available from Joanne, Rachel, and Russet for analysis of vignetting in the focal plane.


ACTION ITEMS:

open from previous meetings:

  • Russet will iterate with the UVA team to clarify some of the language in the TripleSpec User's Guide. Status: OPEN. Still ongoing. Russet needs to narrow down specific questions that are not clear from the manual, then Mike S and John W will help clarify the language.
  • For the NA2 baffle project, Joanne will send a flat field image around to work with. Based on the information from Joanne Nancy will work with John Wilson to add the amplitude of the current loss of signal in the corners of the ARCTIC FOV to the NA2 Baffle Project Description document. Status: OPEN. Joanne was unpacking the flats as we spoke and is preparing to transfer them to Nancy's DropBox account.
  • Ben and Nancy will discuss the ARCTIC flats issue on an upcoming 3.5m engineering telecon to figure out what is acceptable for the Observing Specialists, and come up with a policy regarding how late the observers can keep them to take flats. Status: OPEN. Nancy compiling a policy that might go on users' wiki.
  • UC reps should poll users about which DIS channel is more important to them in the event that we have to choose just one channel to pump an improve the scattered light. Status: CLOSED. GSU, UW answered and said both channels are important to various users.

new from this meeting:

  • All UC reps: Please notify users of constraints on Tspec slit use and forward any interest in using other slits to committee.

Next meeting:

The next Users Committee will be on Monday 11 December at 11:00 am MST. We will discuss the Spring 2018 telecon schedule in December. Nancy also asked if anyone would be at the January AAS meeting; Rachel indicated that she would be there.

Note: See TracWiki for help on using the wiki.